OpenStudy (bobobobo):

Let t(x)=-2x^3+5x-7. Find the new function for t(x) under the following transformations. Express t(x) as a cubic polynomial in the form ax^3+bx^2+cx+d. A. Shift t(x) up 7 units and left 5 units B.Reflect t(x) across the y-axis c. Stretch t(x) horizontally so that it is twice as wide as the original function.

OpenStudy (bobobobo):

please help lol

OpenStudy (welshfella):

to shift the function up 7 units add 7 to t(x) the rule to shift a function a units to the left is f(x) ----> f(x+a) so for t(x) you replace each each x by x + 5 and simplify

OpenStudy (welshfella):

So i'll start you off with A:- adding 7 we get -2x^3 + 5x - 7 + 7 = -2x^3 + 5x Now replace the x by x + 5 and express it in the form required in the question.

OpenStudy (bobobobo):

2(x+5)^3+5(x+5). =2x^3+30x^2+155x+275

OpenStudy (bobobobo):

I just simplified it

OpenStudy (welshfella):

looks good

OpenStudy (bobobobo):

is that the answer?

OpenStudy (welshfella):

thats the answer to A

OpenStudy (welshfella):

B when x is reflected acros the y axis the y values stay the same but the x values change to -x. So if you replace each x by -x in -2x^3 + 5x - 7 you have the answer to B

OpenStudy (welshfella):

* when f(x) is reflected

OpenStudy (welshfella):

- I should have said t(x)!

OpenStudy (bobobobo):

so its just 2x^3-5x-7?

OpenStudy (welshfella):

yeah

OpenStudy (welshfella):

C:- for a horizontal stretch of twice the width you multiply the whole function by 1/2

OpenStudy (bobobobo):

the original function correct?

OpenStudy (welshfella):

OH Hold on I'm not sure about that I'll check that out.

OpenStudy (bobobobo):

oh ok lol

OpenStudy (welshfella):

I was wrong you need to replace the x by 0.5x then simplify

OpenStudy (welshfella):

so we have -2(0.5x)^3 + 5(0.5x) - 7

OpenStudy (bobobobo):

lol my brains dead. been up since 11 at night. ok so -.25^3+2.5x-7?

OpenStudy (welshfella):

lol sorry You are correct

OpenStudy (bobobobo):

oh lol so is that it?

OpenStudy (welshfella):

Yes I have assumed that the 3 transformations are done on the original t(x) . Not consecutive. The question does not make that clear.

OpenStudy (bobobobo):

oh ok thank you so much. you mind helping me out on organizing the quadratic formula?

OpenStudy (welshfella):

ok

OpenStudy (bobobobo):

12x+13-5x^2 I have it as -12+-sqrt12^2-4(-5)13 all over 2(-5)

OpenStudy (welshfella):

that's correct

OpenStudy (bobobobo):

yes however I cannot seem to get the answer.

OpenStudy (bobobobo):

I just want to simplify it

OpenStudy (welshfella):

work out the stuff thats in the brackets first 12^2 - 4*-5*13 = 144 + 260 = 404

OpenStudy (welshfella):

so we have (-12 +/- sqrt 404 ) / -10

OpenStudy (bobobobo):

perhaps you may assist someone else right now. I have to go to the restroom for an emergency lol ill be right back

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\) may u plz close this